if cotxcot7x+cotxcot4x+cot4xcot7x=1 then general solution of x is?

Grade:11

1 Answers

Lab Bhattacharjee
8 years ago
Multiplyboth sides by tanx tan4x tan7x,  tanx+tan4x+tan7x=tanx tan4x tan7x
 
Now tan(A+B+C)=( tanx+tan4x+tan7x-tanx tan4x tan7x)/(1- tanxtan4x-tan7x tan4x-tanx tan7x)
 
So, tan(x+4x+7x)=0, implies x+4x+7x=n pi where n is any integer